# Continuous Urban Change Detection from Satellite Image Time Series with Temporal Feature Refinement and Multi-Task Integration

### v1

@@ IV-B Baseline and benchmark methods
We selected a comprehensive set of baseline and benchmark methods for quantitative and qualitative comparisons with the proposed methods. These selected methods are listed under the categories bi-temporal change detection, multi-temporal change detection, and segmentation methods below.
Bi-temporal urban change detection methods
@@
Multi-temporal urban change detection methods
@@
    L-UNet [26] combines a U-Net for feature extraction in image time series with LSTM modules [28] that obtain rich temporal information from the extracted feature maps.
@@
    Multi-Task L-UNet [27] adds a semantic decoder to L-UNet to segment buildings in the first and last images of a time series.
Building segmentation methods
@@
    U-Net [10] is an encoder-decoder ConvNet widely used for semantic segmentation tasks, including satellite image segmentation.
@@
    HRNet [50], or High-Resolution Net, is a ConvNet that maintains high-resolution representations through the whole network. For semantic segmentation, the high-resolution feature map is combined with the upsampled lower-resolution maps [11].
The Siamese ConvNets Siam-Diff, SNUNet, and DTCDSCN are commonly used as change detection baselines, whereas BIT, AMTNet, and ScratchFormer represent recent methods combining Siamese ConvNets with transformers. On the other hand, L-UNet and Multi-Task L-UNet are benchmark methods for multi-temporal change detection. Finally, HRNet was employed by the winning approach of the SpaceNet 7 challenge for building segmentation, whereas U-Net was used as the baseline method [40]. Both methods also served as a LULC segmentation baseline for the WUSU dataset [39]. Finally, it should be noted that Multi-Task L-UNet and DTCDSCN are multi-task methods that perform change detection and building segmentation.
@@ IV-C Model evaluation
Two accuracy metrics were used for the quantitative assessment of predicted changes: F1 score and intersection over union (IoU). The combination of F1 score and IoU is commonly used for performance assessments in CD studies, e.g., [35, 36]. Formulas for the metrics are given in Eq. 12 and 13, where TP, FP, and FN represent the number of true positive, false positive, and false negative pixels, respectively.
@@
  • Segmentation measures the accuracy of the building predictions corresponding to the first and last image of a time series.
@@ IV-D Implementation details
We implement the proposed method using the deep learning framework PyTorch [52]. To efficiently reshape feature maps, we use the rearrange operation [^1] See https://einops.rocks/api/rearrange/. For the MTI module, we used the pgmpy [^2] See https://pgmpy.org/index.html package to implement the Markov network and perform belief propagation. Models were trained for a maximum duration of 100 epochs on NVIDIA GeForce RTX 3080 graphics cards, using early stopping with patience 10 to prevent models from overfitting to the training set. AdamW was used as optimizer [53] with a linear learning rate scheduler. The remainder of this section describes the training setup in detail.
@@
Multi-Task L-UNet requires additional hyper-parameters to balance the contribution of the segmentation and change loss terms, which we adopted from the paper [27]. All bi-temporal urban change detection methods were trained on the cyclic edges setting (see Figure 5(b)).
@@ V Experiment results
In this section, we present the quantitative and qualitative results on the SpaceNet 7 and WUSU datasets, and the ablation study results. It should be noted that all accuracy values are reported on the respective test sets and correspond to the mean values obtained from five models. These were trained with different seeds using the best hyper-parameters determined with a grid search (see Section IV-D). On the other hand, the median model is used for the qualitative results.
@@ V-A SpaceNet 7
The quantitative results for the SpaceNet 7 dataset are listed in Table I. The proposed method achieved the highest F1 scores and IoU values for both urban change detection tasks (i.e., bi-temporal and continuous). For bi-temporal change detection, methods using multi-temporal information (L-UNet and Multi-Task L-UNet) clearly outperform bi-temporal methods. Among the latter methods, the highest accuracy values were achieved by the ConvNet-based methods Siam-Diff and SNUNet, in addition to the transformer-based method ScratchFromer. For building segmentation, the best performance was also achieved by the proposed method, followed by the segmentation methods U-Net and HRNet and the multi-task methods Multi-Task L-UNet and DTCDSCN.
TABLE I: Quantitative results on the SpaceNet 7 test areas. The best and second-best performance is highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
F1 IoU F1 IoU F1 IoU
Siam-Diff 0.453 0.293 0.273 0.158 - -
SNUNet 0.454 0.294 0.300 0.177 - -
DTCDSCN 0.413 0.260 0.250 0.143 0.488 0.323
BIT 0.386 0.239 0.275 0.160 - -
AMTNet 0.424 0.269 0.282 0.164 - -
ScratchF. 0.469 0.306 0.326 0.195 - -
L-UNet 0.519 0.350 - - - -
MT L-UNet 0.515 0.347 - - 0.512 0.344
U-Net - - - - 0.527 0.359
HRNet - - - - 0.527 0.358
Proposed 0.551 0.381 0.414 0.261 0.596 0.424
Qualitative results for the SpaceNet 7 dataset are shown for urban change detection and building segmentation in Figures 6 and 7, respectively. The change detection results are shown for test sites located in the United States, China, the Philippines, and Australia (top to bottom). Across these sites, the proposed method detects urban changes more accurately than other methods. For example, our method produces the fewest FP pixels (green) for the site located in the Philippines (third row). In addition, it reduces the number of FN pixels (magenta) compared to the other methods for the sites located in China (second row) and Australia (fourth row). The proposed method also accurately maps buildings, as demonstrated in Figure 7 for test sites located (from top to bottom) in the United States (first and second row), the Netherlands (third row), and Australia (fourth row).
Figure 6: Qualitative comparison of the predicted urban changes between the first (first column) and last (second column) image of a time series for a selection of the SpaceNet 7 test areas. Columns three to seven show the benchmarks and the last column shows the proposed method (White: TPs, Black: TNs, Green: FPs, and Magenta: FNs).
Figure 7: Qualitative comparison of the built-up area predictions corresponding to the last image (first column) of a time series for a selection of the SpaceNet 7 test areas. Columns two to four show the benchmarks and the last column shows the proposed method (White: TPs, Black: TNs, Green: FPs, and Magenta: FNs).
@@ V-B WUSU
The quantitative results for the WUSU datasets are listed in Table II. In general, change detection performance on the WUSU dataset is lower than on the SpaceNet 7 dataset. For bi-temporal change detection, only ScratchFormer, L-UNet, and the multi-task methods exceed an F1 score of 0.275 and an IoU value of 0.160. In comparison, accuracy values for the continuous change detection task are slightly higher, except for the proposed method. Overall, our method outperforms all other methods on both change detection tasks. For building segmentation, F1 scores of 0.500 and IoU values of 0.300 are exceeded by all methods with a segmentation output except Multi-Task L-UNet. U-Net and HRNet outperform the multi-task methods; however, the best results were achieved by the proposed method with an F1 score of 0.663 and an IoU value of 0.496.
TABLE II: Quantitative results on the WUSU test areas. The best and second-best performance is highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
F1 IoU F1 IoU F1 IoU
Siam-Diff 0.175 0.096 0.236 0.134 - -
SNUNet 0.188 0.104 0.211 0.118 - -
DTCDSCN 0.278 0.162 0.318 0.189 0.539 0.369
BIT 0.213 0.120 0.314 0.187 - -
AMTNet 0.187 0.104 0.264 0.152 - -
ScratchF. 0.324 0.193 0.353 0.214 - -
L-UNet 0.279 0.162 - - - -
MT L-UNet 0.276 0.161 - - 0.479 0.315
U-Net - - - - 0.576 0.406
HRNet - - - - 0.611 0.440
Proposed 0.440 0.282 0.389 0.242 0.663 0.496
Qualitative results for the WUSU dataset are shown for urban change detection and building segmentation in Figures 8 and 9, respectively. All change detection methods considered in this visual comparison suffer from a large number of FN pixels (purple). In contrast, the proposed method is considerably more effective in detecting urban changes. In addition, the segmentation results for the last image in the time series (see Figure 9) demonstrate that the selected methods accurately map built-up area, even under challenging conditions with considerable off-nadir angles and shadows. Most incorrect predictions (FNs and FPs) are present at the edges of buildings, but a few buildings were also completely omitted. In general, the proposed method produces fewer FPs and FNs than the selected methods in this comparison.
Figure 8: Qualitative comparison of the predicted urban changes between the first (first column) and last (second column) image of a time series for a selection of the WUSU test areas. Columns three to seven show the benchmarks and the last column shows the proposed method (White: TPs, Black: TNs, Green: FPs, and Magenta: FNs).
Figure 9: Qualitative comparison of the built-up area predictions corresponding to the last image (first column) of a time series for a selection of the WUSU test areas. Columns two to four show the benchmarks and the last column shows the proposed method (White: TPs, Black: TNs, Green: FPs, and Magenta: FNs).
@@ V-C Ablation study
Loss function and TFR module: To measure the contribution of applying the loss function to additional change edges in image time series, we perform an ablation experiment that tests different change loss settings on the SpaceNet 7 and WUSU datasets. Furthermore, we investigate the contribution of the TFR module by removing it from the model. The results listed in Table III show that adding change edges in the loss function generally improves performances of the segmentation and change detection tasks. Regarding the TFR module, the results in Table III demonstrate a clear performance gain for both tasks from introducing transformer-based feature refinement. Specifically, adding the TFR module improves segmentation and change detection performance on both datasets considerably.
TABLE III: Ablation results for the loss function and the TFR module on the SpaceNet 7 (SN7) and WUSU datasets. The change loss settings cyclic and dense are equivalent for the WUSU dataset since its time series consists of three images.
Dataset Change loss TFR module Change detection Segmentation
@@
F1 IoU F1 IoU F1 IoU
SN7 adjacent 0.520 0.352 0.384 0.238 0.584 0.413
cyclic 0.532 0.363 0.377 0.233 0.581 0.409
dense 0.537 0.367 0.397 0.248 0.593 0.422
dense 0.490 0.324 0.314 0.186 0.549 0.379
WUSU adjacent 0.356 0.217 0.373 0.229 0.650 0.482
cyclic 0.420 0.266 0.391 0.243 0.660 0.493
cyclic 0.342 0.208 0.364 0.223 0.583 0.412
MTI module: We perform another ablation experiment investigating performance gains from leveraging additional edges in the MTI module. A degenerate Markov network that uses no change information (i.e., only the segmentation information represented as nodes) is added as a baseline. Table V-C lists the ablation results for the SpaceNet 7 and WUSU datasets. Introducing change information in the MTI module results in considerable change detection performance gains compared to the degenerate setting while also providing minor segmentation performance gains. For example, using the adjacent setting improves change detection performance (first-last) by 44.8 % and 61.6 % in terms of F1 score and IoU, respectively, compared to the degenerate setting on the SpaceNet 7 dataset. On the WUSU dataset, the corresponding performance improvements are 57.5 % and 73.6 % for the F1 score and IoU, respectively (cyclic loss scenario). Table V-C also shows that introducing change information beyond adjacent edges results in further change detection performance gains, albeit to a much lesser extent. For example, compared to only using adjacent change information, dense information improved the F1 score and IoU values by 2.0 % and 3.0 %, respectively. In the context of the WUSU dataset, additional change information (i.e., adjacent vs. cyclic) improved the F1 score and IoU values by 1.6 % and 2.2 %, respectively. Therefore, the ablation study demonstrates that the proposed MTI module effectively integrates the outputs of the segmentation and change detection tasks at inference time.
TABLE IV: Ablation results for the MTI module on the SpaceNet 7 and WUSU datasets. The change loss settings cyclic and dense are equivalent for the WUSU dataset since its time series consists of three images.
Dataset Change loss Edges Change detection Segmentation
@@
F1 IoU F1 IoU F1 IoU
SpaceNet 7 adjacent degenerate 0.369 0.226 0.170 0.093 0.584 0.413
adjacent 0.533 0.363 0.398 0.248 0.586 0.415
[…]
adjacent 0.526 0.357 0.391 0.243 0.583 0.411
cyclic 0.535 0.366 0.394 0.246 0.583 0.412
[…]
adjacent 0.540 0.370 0.410 0.258 0.595 0.424
cyclic 0.547 0.377 0.412 0.260 0.596 0.424
dense 0.551 0.381 0.414 0.261 0.596 0.424
WUSU adjacent degenerate 0.244 0.139 0.186 0.103 0.650 0.482
adjacent 0.409 0.257 0.368 0.226 0.656 0.488
[…]
adjacent 0.433 0.276 0.389 0.242 0.663 0.497
cyclic 0.440 0.282 0.389 0.242 0.663 0.496
Time series length: To investigate the effect of satellite image time series length on performance, we tested the proposed network with different settings for $T$. For the change loss and MTI module, the maximum number of edges (i.e., dense edge setting) for a given $T$ was used (see Equation 6). However, it should be noted that for lengths $T=3$ and $T=2$, the edge settings dense are equivalent to cyclic and adjacent, respectively. The results of this experiment (Figure 10) show that adding intermediate images has a positive impact on bi-temporal change detection performance. Specifically, while the proposed method performed worse for $T=2$ compared to SOTA methods (Table I), our model outperforms these methods, including multi-temporal change detection methods, at longer time series lengths. Therefore, it effectively uses multi-temporal information to improve change detection performance.

### v2

@@ IV-B Baseline and benchmark methods
We selected a comprehensive set of baseline and benchmark methods for quantitative and qualitative comparisons with the proposed methods. These selected methods grouped into two categories are listed below.
Bi-temporal change detection methods
@@
Multi-temporal change detection/segmentation methods
@@
    L-UNet [27] employs a shared U-Net for multi-scale feature extraction in SITS and uses LSTM modules [29] for temporal modeling. The LSTM modules produce a single multi-scale feature map which is transformed into the output feature map using a U-Net decoder.
@@
    Multi-Task L-UNet [28] adds a semantic U-Net decoder to L-UNet to segment buildings in the first and last images of a time series.
@@
    U-TAE [33] uses a shared U-Net encoder to extract multi-scale feature maps for the SITS. A temporal attention encoder (L-TAE [47] is then used to collapse the temporal dimension, before using the U-Net decoder to produce a single output feature map.
@@
    TSViT [34] splits a SITS into non-overlapping patches in space and time which are tokenized and subsequently processed by a temporo-spatial encoder. A segmentation head reassembles the encoded features into a single output feature map.
  5. 11.
    U-TempoNet [35] uses a shared ConvNet encoder to extract multi-scale feature maps for all images. Subsequently, a single multi-scale feature map, obtained through temporal modeling with a bidirectional LSTM, is processed with a decoder to produce the output feature map.
The Siamese ConvNets Siam-Diff, SNUNet, and DTCDSCN are commonly used as change detection baselines, whereas BIT, AMTNet, and ScratchFormer represent recent methods combining Siamese ConvNets with transformers. On the other hand, L-UNet and Multi-Task L-UNet are benchmark methods for multi-temporal change detection.It should be noted that Multi-Task L-UNet and DTCDSCN are multi-task methods that perform change detection and building segmentation. Finally, U-TAE, TSViT and U-TempoNet are recent segmentation methods for SITS inputs that can be adopted for multi-temporal change detection without architectural changes.
@@ IV-C Model evaluation
Three accuracy metrics were used for the quantitative assessment of model predictions: F1 score, intersection over union (IoU), and overall accuracy (OA). Formulas for the metrics are given below (Equations 12, 13, and 14), where TP, TN, FP, and FN represent the number of true positive, true negative, false positive, and false negative pixels, respectively.
@@
  • Segmentation measures the accuracy of the building predictions corresponding to the last image of a time series.
@@ IV-D Implementation details
We implement the proposed method using the deep learning framework PyTorch [55]. In addition, the einops package [56] was used to efficiently reshape feature maps, and the pgmpy package [57] to implement the Markov network and perform belief propagation. Models were trained for a maximum duration of 100 epochs on NVIDIA GeForce RTX 3080 graphics cards, using early stopping with patience 10 to prevent models from overfitting to the training set. AdamW was used as optimizer [58] with a linear learning rate scheduler. The remainder of this section describes the training setup in detail.
@@
Multi-Task L-UNet requires additional hyper-parameters to balance the contribution of the segmentation and change loss terms, which we adopted from the paper [28]. All bi-temporal urban change detection methods were trained on the cyclic edges setting (see Fig. 5(b)).
@@ V Experiment results
In this section, we present the quantitative and qualitative results on the SpaceNet 7 and WUSU datasets, and the ablation study results. It should be noted that all accuracy values are reported on the respective test sets and correspond to the mean values obtained from five models. These were trained with different seeds using the best hyper-parameters determined with a grid search (see Section IV-D). On the other hand, the median model is used for the qualitative results, which are only shown for competitive methods selected based on the quantitative results.
@@ V-A SpaceNet 7
The quantitative results for the SpaceNet 7 dataset are listed in Table I. The proposed method achieved the highest F1 scores and IoU values for both urban change detection tasks (i.e., bi-temporal and continuous). Several multi-temporal models (L-UNet, Multi-Task L-UNet, and U-TempoNet) outperform bi-temporal change detection methods on the bi-temporal task while others are less effective (U-TAE and TSViT). Among the bi-temporal methods, ScartchFormer and the ConvNet-based methods Siam-Diff and SNUNet achieved the highest accuracy values. For building segmentation, the proposed method also outperformed the other multi-task methods including DTCDSCN and Multi-Task L-UNet.
TABLE I: Quantitative results on the SpaceNet 7 test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
(F1 / IoU / OA) (F1 / IoU / OA) (F1 / IoU / OA)
Siam-Diff 0.453 / 0.293 / 98.8 0.273 / 0.158 / 99.5 -
SNUNet 0.454 / 0.294 / 98.8 0.300 / 0.177 / 99.6 -
DTCDSCN 0.413 / 0.260 / 98.7 0.250 / 0.143 / 99.6 0.488 / 0.323 / 92.3
BIT 0.386 / 0.239 / 99.0 0.275 / 0.160 / 99.6 -
AMTNet 0.424 / 0.269 / 98.7 0.282 / 0.164 / 99.6 -
ScratchFormer 0.468 / 0.305 / 98.9 0.328 / 0.196 / 99.6 -
L-UNet 0.519 / 0.350 / 98.9 - -
MT L-UNet 0.515 / 0.347 / 98.9 - 0.512 / 0.344 / 92.7
U-TAE 0.366 / 0.225 / 97.5 - -
TSViT 0.168 / 0.092 / 97.2 - -
U-TempoNet 0.494 / 0.328 / 98.8 - -
Proposed 0.551 / 0.381 / 99.0 0.414 / 0.261 / 99.7 0.596 / 0.424 / 94.3
Fig. 6 and 7 show qualitative change detection and building segmentation results for two SpaceNet 7 test sites located in Australia and the United States, respectively. For both sites, the proposed method detects urban changes more accurately than competing methods (DTCDSCN, ScratchFormer, and Multi-Task L-UNet). In particular, the continuous change detection results for consecutive image pairs (rows two to five) show a better agreement with the label than those of the other methods. In addition, the change outputs of the proposed method show a high level of consistency, meaning that the aggregated continuous changes correspond to the bi-temporal changes between the first and the last image (top row). Finally, the proposed method maps buildings with more detail than the competing methods, as shown in the bottom row of Fig. 6 and 7.
Figure 6: Qualitative urban change and building segmentation results for a SpaceNet 7 test site located in Australia. The PlanetScope satellite image time series is shown on the left, and the model outputs and the label are shown on the right. The top row shows the changes between the first and last image of the time series, rows two to five show the continuous changes between consecutive image pairs, and the bottom row the buildings segmentation corresponding to the last image.
Figure 7: Qualitative urban change and building segmentation results for a SpaceNet 7 test site located in the United States. The PlanetScope satellite image time series is shown on the left, and the model outputs and the label are shown on the right. The top row shows the changes between the first and last image of the time series, rows two to five show the continuous changes between consecutive image pairs, and the bottom row the buildings segmentation corresponding to the last image.
@@ V-B WUSU
The quantitative results for the WUSU datasets are listed in Table II. In general, change detection performance on the WUSU dataset is lower than on the SpaceNet 7 dataset. For bi-temporal change detection, only ScratchFormer, L-UNet, and Multi-Task L-UNet exceed an F1 score of 0.275 and an IoU value of 0.160. In comparison, accuracy values for the continuous change detection task are slightly higher, except for the proposed method. Overall, our method outperforms all other methods on both change detection tasks. This also applied to the building segmentation task, where the proposed method achieved an F1 score of 0.663 and an IoU value of 0.496.
TABLE II: Quantitative results on the WUSU test areas. The best and second-best performances are highlighted in red and blue, respectively. ”-” denotes that the accuracy metric does not apply to a specific method since the corresponding variable is not predicted.
@@
(F1 / IoU / OA) (F1 / IoU / OA) (F1 / IoU / OA)
Siam-Diff 0.175 / 0.096 / 96.1 0.236 / 0.134 / 97.4 -
SNUNet 0.188 / 0.104 / 94.9 0.211 / 0.118 / 96.5 -
DTCDSCN 0.278 / 0.162 / 96.2 0.318 / 0.189 / 97.6 0.539 / 0.369 / 84.2
BIT 0.213 / 0.120 / 96.5 0.314 / 0.187 / 97.9 -
AMTNet 0.187 / 0.104 / 95.9 0.264 / 0.152 / 97.6 -
ScratchFormer 0.324 / 0.193 / 96.6 0.352 / 0.214 / 97.8 -
L-UNet 0.279 / 0.162 / 96.0 - -
MT L-UNet 0.276 / 0.161 / 96.1 - 0.479 / 0.315 / 83.1
U-TAE 0.267 / 0.154 / 92.1 - -
TSViT 0.219 / 0.123 / 92.1 - -
U-TempoNet 0.246 / 0.141 / 95.0 - -
Proposed 0.440 / 0.282 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.496 / 88.9
Fig. 8 and 9 show qualitative change detection and building segmentation results for two WUSU test sites located in Wuhan’s Jiang’an District in China. The Gaofen-2 SITS for both sites highlight the challenging nature of the WUSU dataset. In particular, the images were acquired under different atmospheric conditions and they have large off-nadir angles and contain shadows. Despite that, the change detection results produced by the proposed method show good agreement with the label, especially in comparison with the competing methods. As for SpaceNet 7, the proposed method achieved a high level of consistency between the continuous change detection outputs (rows two and three) and the change output between the first and last image (top row). In contrast, the bi-temporal change detection methods identified changes between the first and last images (ch $t1\rightleftharpoons t3$) that are present in neither of the continuous change rows (ch $t1\rightleftharpoons t2$ and ch $t2\rightleftharpoons t3$). Finally, the bottom row demonstrates that all methods accurately map buildings, but the proposed method achieved more detailed building delineations than its competitors. It should also be noted that the label does not distinguish individual buildings in very dense built-up areas (e.g., the bottom left area in Fig. 8).
Figure 8: Qualitative urban change and building segmentation results for a WUSU test site located in Wuhan’s Jiang’an District, China. The Gaofen-2 satellite image time series is shown on the left, and the model outputs and the label are shown on the right. The top row shows the changes between the first and last image of the time series, rows two and three show the continuous changes between consecutive image pairs, and the bottom row the buildings segmentation corresponding to the last image.
Figure 9: Qualitative urban change and building segmentation results for a WUSU test site located in Wuhan’s Jiang’an District, China. The Gaofen-2 satellite image time series is shown on the left, and the model outputs and the label are shown on the right. The top row shows the changes between the first and last image of the time series, rows two and three show the continuous changes between consecutive image pairs, and the bottom row the buildings segmentation corresponding to the last image.
@@ V-C Ablation study
Loss function: Table III shows the ablation results for different change loss edge settings (first-last, adjacent, cyclic, and dense) and the segmentation loss. It should be noted that the settings adjacent, cyclic, and dense all require building annotations for each image in the time series as labels (see Section III-F, whereas first-last only requires building annotations for the first and last images. In general, considering additional change edges in the loss function improves continuous change detection performance. However, a single change loss term suffices for the bi-temporal change detection task. The segmentation loss term improves performance for both change detection tasks, which holds for all change loss settings. Therefore, the optimal loss setting consists of a change loss with dense edges combined with a segmentation loss.
TABLE III: Ablation results for the loss function with different edge settings for the change loss term. The best and second-best performances are highlighted in red and blue, respectively.
Dataset Change loss Seg loss Change detection Segmentation
@@
(F1 / IoU / OA) (F1 / IoU / OA) (F1 / IoU / OA)
SN7 first-last 0.453 / 0.294 / 98.7 0.253 / 0.145 / 99.6 -
first-last 0.519 / 0.350 / 98.8 0.323 / 0.193 / 99.7 0.565 / 0.394 / 93.8
adjacent 0.479 / 0.315 / 98.8 0.357 / 0.217 / 99.6 -
adjacent 0.520 / 0.352 / 98.9 0.384 / 0.238 / 99.6 0.584 / 0.413 / 94.1
cyclic 0.516 / 0.348 / 98.8 0.361 / 0.220 / 99.6 -
cyclic 0.532 / 0.363 / 98.9 0.377 / 0.233 / 99.7 0.581 / 0.409 / 94.0
dense 0.511 / 0.343 / 98.8 0.364 / 0.223 / 99.6 -
dense 0.537 / 0.367 / 98.8 0.397 / 0.248 / 99.7 0.593 / 0.422 / 94.3
WUSU first-last 0.274 / 0.159 / 95.7 0.234 / 0.133 / 96.3 -
first-last 0.401 / 0.251 / 96.5 0.297 / 0.175 / 98.3 0.650 / 0.482 / 88.4
adjacent 0.251 / 0.144 / 96.4 0.297 / 0.175 / 97.9 -
adjacent 0.356 / 0.217 / 95.3 0.373 / 0.229 / 98.2 0.650 / 0.482 / 88.6
cyclic 0.328 / 0.196 / 96.0 0.289 / 0.169 / 98.2 -
cyclic 0.420 / 0.266 / 96.6 0.391 / 0.243 / 98.3 0.660 / 0.493 / 88.8
TFR module: We perform an additional ablation experiment investigating the contribution of the TFR module and testing it with different recurrent sequence models, including recurrent neural networks (RNNs) [60] and LSTMs [29]. Furthermore, we run all settings with and without the MTI module to highlight the complementary nature of the modules. Table IV shows that adding a temporal sequence module improves performances across all three tasks. Among the sequence modules, the transformer encoder is most effective for the SpaceNet 7 dataset. This also holds for the segmentation task on the WUSU dataset. On the other hand, all sequence models achieved similar performances on the WUSU change detection tasks.
TABLE IV: Ablation results for the TFR module with different sequence models on the SpaceNet 7 (SN7) and WUSU datasets. The maximum number of edges was used for each experiment in the change loss. The best and second-best performances are highlighted in red and blue, respectively.
Dataset TFR module Seq. model Change detection Segmentation
@@
(F1 / IoU / OA) (F1 / IoU / OA) (F1 / IoU / OA)
SN7 - 0.511 / 0.343 / 98.9 0.344 / 0.208 / 99.6 0.576 / 0.404 / 94.2
RNN 0.525 / 0.356 / 98.9 0.377 / 0.232 / 99.6 0.572 / 0.401 / 93.8
LSTM 0.547 / 0.377 / 99.0 0.402 / 0.252 / 99.6 0.588 / 0.417 / 94.4
TE 0.551 / 0.381 / 99.0 0.414 / 0.261 / 99.7 0.596 / 0.424 / 94.3
WUSU - 0.392 / 0.245 / 96.6 0.339 / 0.205 / 97.6 0.649 / 0.480 / 88.4
RNN 0.440 / 0.282 / 96.9 0.390 / 0.243 / 98.2 0.658 / 0.491 / 88.4
LSTM 0.441 / 0.283 / 96.9 0.387 / 0.240 / 98.2 0.658 / 0.490 / 88.5
TE 0.440 / 0.282 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.496 / 88.9
MTI module: We perform a third ablation experiment investigating performance gains from leveraging additional edges in the MTI module. A degenerate Markov network that uses no change information (i.e., only the segmentation information represented as nodes) is added as a baseline. Table V shows that introducing change information in the MTI module results in considerable change detection performance gains compared to the degenerate setting, as well as minor segmentation performance gains. For example, using the adjacent setting improves change detection performance (first-last) by 44.8 % and 61.6 % in terms of F1 score and IoU, respectively, compared to the degenerate setting on the SpaceNet 7 dataset. On the WUSU dataset, the corresponding performance improvements are 57.5 % and 73.6 % for the F1 score and IoU, respectively (cyclic loss scenario). Table V also shows that introducing change information beyond adjacent edges results in further change detection performance gains, albeit to a much lesser extent. For example, compared to only using adjacent change information, dense information improved the F1 score and IoU values by 2.0 % and 3.0 %, respectively. In the context of the WUSU dataset, additional change information (i.e., adjacent vs. cyclic) improved the F1 score and IoU values by 1.6 % and 2.2 %, respectively. Therefore, the ablation study demonstrates that the proposed MTI module effectively integrates the outputs of the segmentation and change detection tasks at inference time.
TABLE V: Ablation results for the MTI module on the SpaceNet 7 (SN7) and WUSU datasets. The edge settings cyclic and dense are equivalent for the WUSU dataset since its time series consists of three images. The best and second-best performances are highlighted in red and blue, respectively.
Dataset MTI module Edge setting Change detection Segmentation
Bi-temporal Continuous
(F1 / IoU / OA) (F1 / IoU / OA) (F1 / IoU / OA)
SN7 - 0.537 / 0.367 / 98.8 0.397 / 0.248 / 99.7 0.593 / 0.422 / 94.3
degenerate 0.373 / 0.229 / 97.8 0.173 / 0.095 / 98.8 0.593 / 0.422 / 94.3
adjacent 0.540 / 0.370 / 99.0 0.410 / 0.258 / 99.7 0.595 / 0.424 / 94.4
cyclic 0.547 / 0.377 / 99.0 0.412 / 0.260 / 99.7 0.596 / 0.424 / 94.3
dense 0.551 / 0.381 / 99.0 0.414 / 0.261 / 99.7 0.596 / 0.424 / 94.3
WUSU - 0.420 / 0.266 / 96.6 0.391 / 0.243 / 98.3 0.660 / 0.493 / 88.8
degenerate 0.275 / 0.159 / 93.8 0.203 / 0.113 / 95.5 0.660 / 0.493 / 88.8
adjacent 0.433 / 0.276 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.497 / 88.9
cyclic 0.440 / 0.282 / 97.0 0.389 / 0.242 / 98.3 0.663 / 0.496 / 88.9
Time series length: To investigate the effect of SITS length on performance, we tested the proposed network with different settings for $T$. For the change loss and MTI module, the maximum number of edges (i.e., dense edge setting) for a given $T$ was used (see Equation 6). However, it should be noted that for lengths $T=3$ and $T=2$, the edge settings dense are equivalent to cyclic and adjacent, respectively. The results of this experiment (Fig. 10) show that adding intermediate images has a positive impact on bi-temporal change detection performance. Specifically, while the proposed method performed worse for $T=2$ compared to SOTA methods (Table I), our model outperforms these methods, including multi-temporal change detection methods, at longer time series lengths. Therefore, it effectively uses multi-temporal information to improve change detection performance.